# A Quantitative Comparison of Pre-Trained Model Registries to Traditional Software Package Registries

### v1

@@ 4.1. Methods
4.1.1. Pilot Study
We define the scope of our review by conducting a pilot study. In the pilot study, we first search for papers about “pre-trained model reuse” using Google Scholar and looked at the first three results, which were (Jiang et al., 2022, 2023b; Gong et al., 2023a). The papers indicated that Hugging Face is the only “open” model registry (Jiang et al., 2022) and is the most popular model registry. Additionally, Hugging Face hosts the largest number of PTM packages, and provides useful tools to facilitate PTM reuse. We then decided to scope down our study on Hugging Face model registry specifically to represent the PTM supply chain, as indicated by Jiang et al. (Jiang et al., 2024b, 2023b).
4.1.2. Search Strategy and Query
The goal of our search is to identify papers that are relevant to the categories of PTM reuse, the PTM supply chain, or the PTM ecosystem. Informed by our pilot study ( section 4.1.1), the final search query we used is indicated in Figure 4. These search queries gave us 45 papers. We then removed duplicate entries, reducing the number of papers to 31.
To verify the efficacy of our search queries, we employed papers from the pilot study as benchmarks to assess each query’s retrieval effectiveness. We ensured that all papers identified in the pilot study were also retrieved by our final search query. This step was essential to confirm the robustness of our search strategy, ensuring it was capable of capturing the most relevant studies. Such a comprehensive approach allowed for an exhaustive review of the literature concerning PTM reuse within its operational ecosystem.
4.1.3. Selection Criteria
The goal of our selection criteria is to identify the most relevant and rigorously supported research that specifically addresses the context and impact of PTM reuse. During our study, we applied two types of criteria: inclusive and exclusive.
Inclusion Criteria We applied the following inclusion criterion: a paper is included if it describes the reuse of models within a specific PTM registry. This criterion excluded papers that apply PTMs to specific tasks, reducing our set from 31 papers to 20.
Exclusion Criteria Our exclusion criterion was that we excluded non-primary sources. We exclude works whose claims are not substantiated directly through qualitative or quantitative methods. This reduced our set from 20 papers to 18.
@@ 4.1.3. Selection Criteria
Motivation claims Articulates the rationale behind a paper’s problem statement, illustrating why the work is significant and worthy of investigation. The reuse of pre-trained models introduces large costs and additional problems of checking whether arbitrary pre-trained models are suitable for the task-specific reuse or not.” (Gong et al., 2023a)
@@
Work claims Refers to the assertions a paper makes based on its collected data and analyses. Hugging Face’s popularity has exponentially increased over time, which is evident from the upward trends in the number of new models, likes, commits, unique authors, and discussions aggregated monthly.” (Castaño et al., 2023)
Engineers follow specific naming practices and encounter challenges that are specific to PTM naming.” (Jiang et al., 2024a)
4.1.4. Data Extraction
@@ 4.2. Analysis and Results
The primary aim of our SLR is to summarize the existing claims about package reuse on Hugging Face and to extract quantifiable measurements from these claims. From the consolidated set of 12 claims, we categorized the basis of each claim into one of five methods: small- and large-scale quantitative measurements, qualitative surveys, interviews, and case studies. Small-scale measurements involved less than 10% of a population, whereas large-scale measurements encompassed more than 10%. After this classification, five quantitatively unevaluated claims remained: one concerning design trends, two regarding selection considerations, and two about documentation and understanding.
The categorization result and extracted themes are detailed in Table 3. Our results provide a thorough knowledge synthesis which are then used to answer RQ2 (section 5).
@@
Trends in Design A small group of contributors owns popular models. Large-Scale Measurement (Castaño et al., 2023)
The Transformers library increases the accessibility of PTM creation and downstream reuse. Large-Scale Measurement, Case Study (Kathikar et al., 2023; Wolf et al., 2020; Jiang et al., 2024b; Castaño et al., 2023)
@@
Selection Considerations $\dagger$ (4) DL-specific attributes such as model architecture, performance, reproducibility, and portability affect PTM selection and reuse. Interviews (Jiang et al., 2023b)
$\dagger$ (2) The traditional attribute of Popularity affects PTM selection and reuse more than Maintenance and Quality. Interviews (Jiang et al., 2023b)
Repository Lifecycle and Maintenance Models receive perfective maintenance over time, with high-maintenance models tending to be more popular, larger, and better documented. Large-Scale Measurement (Castaño et al., 2023)
Documentation and Understanding Model properties are under-documented across Hugging Face. Small-Scale and Large-Scale Measurement, Survey (Kathikar et al., 2023; Taraghi et al., 2024; Montes et al., 2022; Gong et al., 2023a)
@@
Naming models is inconsistent and can inadequately represent model architectures. Large-Scale Measurement, Survey (Jiang et al., 2024a)
@@
Downstream Usage Hugging Face is an exponentially growing platform. Large-Scale Measurement (Castaño et al., 2023)
5. RQ2: Do the qualitative claims about package reuse (PTM) on Hugging Face hold up when quantified?
Finding 4. Our study shows that the Transformers library is preferred in over 80% of PTM descendants, surpassing PyTorch. The rise of the SafeTensors library underscores a shift toward prioritizing security in PTM development.
Finding 5. Figure 6 reveals that Hugging Face has a significantly higher package turnover rate than traditional software registries indicative of a fast-paced, innovation-driven PTM ecosystem. Finding 6. There is a correlation between model popularity and descendant count, indicating that while popular models have more descendants, other factors influence model selection decisions. Finding 7. There is a strong correlation between documentation quality and PTM popularity, with the top 1000 models significantly outperforming the bottom 1000 in documentation, highlighting its importance in model selection.
@@ 5.1. Metrics Developed from Claims
In this section, we explain how we developed quantifiable metrics from the claims identified in our SLR (section 4). The metrics were specifically designed to quantify the hypothesized results inferred from these claims. Drawing on traditional software engineering practices, we adopted metrics that have been previously used to evaluate similar claims in other contexts if available.
Particular attention was paid to metrics that are widely recognized and have been frequently cited in the literature, as well as those that have been implemented across various traditional software registries. This approach ensures that our metrics are grounded in established methodologies and are robust enough to provide meaningful insights.
Table 4 shows the relationship between the claims extracted and the corresponding metrics, along with the expected measurements. This mapping both validaties the claims and contextualizes their implications within the context of package reuse on Hugging Face.
@@ 5.2. Available Datasets
5.2.1. PTM Datasets
In the PTM literature, there are four datasets available publicly: HF Model Metadata (Van Strien, 2022), PTMTorrent (Jiang et al., 2023c), HFCommunity (Ait et al., 2023), and PeaTMOSS (Jiang et al., 2024b).
@@ 5.2.1. PTM Datasets
    PTMTorrent encompasses a snapshot of five model hubs, totaling 15,913 PTM packages as of 08/2023, all formatted in a uniform data schema to facilitate cross-hub mining (Jiang et al., 2023c).
@@
    PeaTMOSS offers comprehensive metadata on 281,638 PTM packages as of 10/2023, including 281,276 from Hugging Face and 362 from PyTorch Hub, along with details on 28,575 GitHub projects that use PTMs as dependencies and 44,337 links from these GitHub repositories back to the PTMs they depend on (Jiang et al., 2024b).
@@
5.2.2. Traditional package datasets
To directly compare measurements between the PTM registry and traditional Software Package Registries (SPRs), we utilized the Ecosyste.ms software package dataset for traditional packages, following the approach outlined in prior work (Schorlemmer et al., 2024). This dataset provides a set of free and open resources for those working to sustain and secure open source software. Ecosyste.ms publishes open data and APIs that map software interdependencies, as well as providing data on the usage, creation, and potential impact of packages.
In this work, we used the version of the dataset from October 2023, as this is the closest in time to when the PeaTMOSS dataset was created. This includes the usage of data from two software package registries: PyPI and NPM.
5.3. $C_{1}$: The Transformers library increases the accessibility of PTM creation and downstream reuse.
@@ 5.2.2. Traditional package datasets
5.3.1. Method
We present the metric we developed for claim 1. Metric 1: Preservation rate of libraries to descendents: Some models support multiple libraries. If the claim holds, then descendants of those models should make use of (and thus support) the more reuse-friendly libraries. We consider each library $L$ in turn to assess how frequently its descendants continue to use it. The ingredients of our measure are: the library $L$ being assessed, the set of base models $B$ that use library $L$ and at least one other library, the set $D_{b}$ of direct descendants of base model $b\in B$, and a function $S(d,L)$ that returns 1 if descendant $d\in D_{b}$ supports library $L$, otherwise 0. We can then calculate the preservation rate $P_{L}$ of a library $L$ as:
@@ 5.3.1. Method
The numerator is the total descendants supporting library $L$, calculated as: $\sum_{b\in B}\sum_{d\in D_{b}}S(d,L)$. The denominator is the total count of direct descendants across all base models in set $B$: $\sum_{b\in B}|D_{b}|$.
This equation will give the percentage of direct descendants that support library $L$. The raw count of descendants supporting the library can be found using the numerator.
5.3.2. Results
Metric 1: Preservation rate of libraries to descendents: As depicted in Figure 5, the Transformers library has the highest survival rate from a parent model to its descendant among all libraries used. Specifically, over 80% of descendant models continue to employ the Transformers library, establishing it as the preferred choice for generating descendant models. Transformers is also the most popular library (has the largest population of models that use it). Contrarily, despite previous studies suggesting high prevalence, PyTorch was not favored among PTM descendants, a departure from claims within our systematic literature review that both Transformers and PyTorch are dominant on Hugging Face (Castaño et al., 2023).
Figure 5. The usage proportion of the top-10 libraries that PTMs utilizing at least two different libraries use on Hugging Face. For PTM packages that leverage support at least two libraries, most packages support the transformers library, followed by the Hugging Face promoted SafeTensors library. Most other libraries have little usage in comparison. In contrast to previous work (Castaño et al., 2023), PyTorch is not one of the most popular library to be supported when a PTM package supports more than one library.
Instead, our findings highlight the Transformers and SafeTensors libraries as the most prevalent, suggesting a community shift towards prioritizing security, as SafeTensors is designed to replace the commonly used Python pickle library with a more secure container for deploying models. We attribute this discrepancy to prior studies’ broader approach, which did not differentiate between model types and analyzed only a single snapshot of Hugging Face, thus lacking the detailed analysis presented here. Our results suggest that PTM synthesizers are willing to compromise on functionality and portability to ensure the distribution of more secure PTM packages.
@@ 5.3.2. Results
5.4.1. Methods
We present the metrics we developed for claim 2. Metric 2: Turnover of Top PTMs: Drawing on prior work characterizing the stability of top packages over time (Dey and Mockus, 2018), we measured the top-$K$ turnover for each registry. Let $S_{current}$ be the set of $K$ most popular packages in the current snapshot, and $S_{last}$ be the set of top $K$ packages in the last snapshot. We also consider $S_{history}$, the set of packages in any snapshot before $S_{last}$. We then distinguish three categories for packages in $S_{current}$:
@@ 5.4.1. Methods
For the measurement, we defined popularity by the number of downloads, and examined the top-1000 packages. We obtained snapshots across four dates for both traditional software and PTMs. Data came from several HuggingFace snapshots (datasets: HF Model Metadata–June 2022, PTMTorrent–May 2023, PeaTMOSS–Oct. 2023). We took a current snapshot of the top 1000 PTMs directly from Hugging Face (April 2024). We used the Ecosyste.ms dataset (NPM, PyPI) for a comparison with traditional software package registries.
@@
5.4.2. Results
Metric 2: Turnover of Traditional Software and PTM registries: Figure 6 shows the results for Hugging Face. The Hugging Face data does not match our interpretation of the claim. About half of the top-1K Hugging Face PTMs turned over in each snapshot. This high turnover rate suggests that packages on Hugging Face have a shorter lifespan, indicating a dynamic PTM environment where the requirements and preferred models frequently change (Gopalakrishna et al., 2023). Further investigation could analyze the lifecycle of these PTMs to determine whether they are newer models briefly appearing or established ones losing prominence, as well as identifying the traits of PTMs or maintainers who consistently stay within the top rankings. This result is consistent with the claim, but implies that popularity is likely driven by performance (a latent variable not present in the claim).
Figure 6 also shows NPM for comparison. The results for PyPI were similar. Traditional registries show stability with their most popular packages. Quantitatively, 2535 distinct packages were in the HuggingFace top-1K, while only 1127 were there for NPM.
@@ 5.4.2. Results
The rapid turnover on Hugging Face compared to the stability observed in traditional software packages indicates that PTM needs are continually evolving, unlike the more established needs in traditional software domains. This evolution often leads to older PTMs being supplanted by newer models that better meet current requirements or offer superior performance, highlighting a market driven by innovation and rapid adaptation. This trend suggests that PTM users are quick to adopt new advancements, reflecting the field’s fast-paced development and the shifting demands across its various application domains.
Metric 3: Number of Descendents of Top Packages: Figure 7 shows the results for this metric. We observe a weak positive correlation between popularity (downloads) and the number of descendent models. Popular models tend to have more descendants, but the correlation was weaker than expected. Notably, some highly popular models had relatively few descendants, while others with similar popularity levels had many. This suggests that while popularity influences the likelihood of a model being chosen for further development, it is not the sole factor. Further analysis, including a breakdown by model task and domain, is necessary to understand if variations in descendant counts are influenced by the specific popularity within less prominent domains or tasks. This analysis could expose the decision-making processes engineers use when selecting models to fine-tune and develop further.

### v2

@@ 4.1. Methods
4.1.1. Pilot Study.
We define the scope of our review by conducting a pilot study. In the pilot study, we first search for papers about “pre-trained model reuse” using Google Scholar and looked at the first three results, which were (Jiang et al., 2022, 2023a; Gong et al., 2023). The papers indicated that Hugging Face is the only “open” model registry (Jiang et al., 2022) and is the most popular model registry. Additionally, Hugging Face hosts the largest number of PTM packages, and provides useful tools to facilitate PTM reuse. We then decided to scope down our study on Hugging Face model registry specifically to represent the PTM supply chain, as indicated by Jiang et al. (Jiang et al., 2024c, 2023a).
4.1.2. Search Strategy and Query.
The goal of our search is to identify papers that are relevant to the categories of PTM reuse, the PTM supply chain, or the PTM ecosystem.
Informed by our pilot study ( section 4.1.1), the final search query we used is indicated in Figure 4. These search queries gave us 45 papers. We then removed duplicate entries, reducing the number of papers to 31. To verify the efficacy of our search queries, we employed papers from the pilot study as benchmarks to assess each query’s retrieval effectiveness. We ensured that all papers identified in the pilot study were also retrieved by our final search query. This step was essential to confirm the robustness of our search strategy, ensuring it was capable of capturing the most relevant studies. Such a comprehensive approach allowed for an exhaustive review of the literature concerning PTM reuse within its operational ecosystem.
4.1.3. Selection Criteria.
The goal of our selection criteria is to identify the most relevant and rigorously supported research that specifically addresses the context and impact of PTM reuse. We developed and applied a set of clearly articulated criteria to ensure consistency and reproducibility in our selection process.
Inclusion Criteria We applied the following inclusion criteria to our systematic literature review: The paper must describe the reuse of models within a specific PTM registry, particularly Hugging Face. Additionally, the study should focus on PTM reuse practices, challenges, or impacts. To ensure academic rigor, we only included papers published in peer-reviewed journals or conference proceedings. Lastly, to maintain relevance to current practices, we limited our search to publications within the last five years (2019-2024).
Exclusion Criteria We applied the following exclusion criteria in our systematic literature review: Papers that only apply PTMs to specific tasks without discussing reuse practices were excluded. We also omitted non-primary sources, such as literature reviews or opinion pieces, to focus on original research. Additionally, we excluded studies whose claims were not substantiated directly through qualitative or quantitative methods, ensuring that our review was based on well-supported findings.
Application of these criteria reduced our initial set of 31 papers to 18 high-quality, relevant studies for in-depth analysis.
@@
Motivation claims Articulates the rationale behind a paper’s problem statement, illustrating why the work is significant and worthy of investigation. The reuse of pre-trained models introduces large costs and additional problems of checking whether arbitrary pre-trained models are suitable for the task-specific reuse or not.” (Gong et al., 2023)
@@
Work claims Refers to the assertions a paper makes based on its collected data and analyses. Hugging Face’s popularity has exponentially increased over time, which is evident from the upward trends in the number of new models, likes, commits, unique authors, and discussions aggregated monthly.” (Castaño et al., 2024)
Engineers follow specific naming practices and encounter challenges that are specific to PTM naming.” (Jiang et al., 2024b)
4.1.4. Data Extraction.
@@ 4.2. Analysis and Results
The primary aim of our SLR is to summarize the existing claims about package reuse on Hugging Face and to extract quantifiable measurements from these claims. From the consolidated set of 12 claims, we categorized the basis of each claim into one of five methods: small- and large-scale quantitative measurements, qualitative surveys, interviews, and case studies. Small-scale measurements involved less than 10% of the population, while large-scale measurements covered more than 10%. After this classification, five quantitatively unevaluated claims remained: one concerning design trends, two regarding selection considerations, and two about documentation and understanding. The categorization result and extracted themes are detailed in Table 3.
Our comprehensive knowledge synthesis addresses RQ2 (section 5), revealing several key insights into PTM reuse research. We observe a significant shift towards large-scale, data-driven analysis, complemented by diverse research methods that collectively strengthen the field’s knowledge base. The consolidation of 30 initial claims into 12 distinct work claims indicates both a convergence of findings and a need for validation through replication studies. Notably, we identified five quantitatively unevaluated claims, highlighting critical areas for future empirical investigation, particularly in design trends, selection considerations, and documentation practices.
Our findings have significant implications for PTM reuse research and practice. For researchers, we provide a roadmap highlighting areas requiring rigorous quantitative evaluation. Practitioners can leverage our identified themes to inform PTM selection and implementation decisions. The emphasis on documentation-related claims underscores its critical importance in the PTM ecosystem for both groups. This comprehensive analysis lays the foundation for targeted, impactful research, contributing to the maturation of the rapidly evolving PTM reuse field.
@@
Trends in Design A small group of contributors owns popular models. Large-Scale Measurement (Castaño et al., 2024)
The Transformers library increases the accessibility of PTM creation and downstream reuse. Large-Scale Measurement, Case Study (Kathikar et al., 2023; Wolf et al., 2020; Jiang et al., 2024c; Castaño et al., 2024)
@@
Selection Considerations $\dagger$ (4) DL-specific attributes such as model architecture, performance, reproducibility, and portability affect PTM selection and reuse. Interviews (Jiang et al., 2023a)
$\dagger$ (2) The traditional attribute of Popularity affects PTM selection and reuse more than Maintenance and Quality. Interviews (Jiang et al., 2023a)
Repository Lifecycle and Maintenance Models receive perfective maintenance over time, with high-maintenance models tending to be more popular, larger, and better documented. Large-Scale Measurement (Castaño et al., 2024)
Documentation and Understanding Model properties are under-documented across Hugging Face. Small-Scale and Large-Scale Measurement, Survey (Kathikar et al., 2023; Taraghi et al., 2024; Montes et al., 2022; Gong et al., 2023)
@@
Naming models is inconsistent and can inadequately represent model architectures. Large-Scale Measurement, Survey (Jiang et al., 2024b)
@@
Downstream Usage Hugging Face is an exponentially growing platform. Large-Scale Measurement (Castaño et al., 2024)
5. RQ2: Qualitative Claim Validation
Finding 4. Our study shows that the Transformers library is preferred in over 80% of PTM descendants, surpassing PyTorch. The rise of the SafeTensors library underscores a shift toward prioritizing security in PTM development. Finding 5. Figure 6 reveals that Hugging Face has a significantly higher package turnover rate than traditional software registries indicative of a fast-paced, innovation-driven PTM ecosystem. Finding 6. Model popularity correlates with descendant count, indicating that while popular models have more descendants, other factors influence model selection decisions. Finding 7. There is a strong correlation between documentation quality and PTM popularity, with the top 1000 models significantly outperforming the bottom 1000 in documentation, highlighting its importance in model selection.
@@ 5.1. Metrics Developed from Claims
In this section, we develop quantifiable metrics from the claims identified in our SLR (section 4). These metrics are designed to test hypothesized results inferred from the claims, drawing on established software engineering practices and previously used metrics where applicable. We prioritized widely recognized, frequently cited metrics that have been implemented across various traditional software registries. This approach grounds our metrics in established methodologies, ensuring robust and meaningful insights. Table 4 shows the relationship between extracted claims, corresponding metrics, and expected measurements, validating the claims and contextualizing their implications for package reuse on Hugging Face.
@@ 5.2. Available Datasets
5.2.1. PTM Datasets.
In the PTM literature, there are four datasets available publicly: HF Model Metadata (Van Strien, 2022), PTMTorrent (Jiang et al., 2023b), HFCommunity (Ait et al., 2023), and PeaTMOSS (Jiang et al., 2024c).
@@
    PTMTorrent encompasses a snapshot of five model hubs, totaling 15,913 PTM packages as of 08/2023, all formatted in a uniform data schema to facilitate cross-hub mining (Jiang et al., 2023b).
@@
    PeaTMOSS offers comprehensive metadata on 281,638 PTM packages as of 10/2023, including 281,276 from Hugging Face and 362 from PyTorch Hub, along with details on 28,575 GitHub projects that use PTMs as dependencies and 44,337 links from these GitHub repositories back to the PTMs they depend on (Jiang et al., 2024c).
@@
5.2.2. Traditional package datasets.
To directly compare measurements between the PTM registry and traditional Software Package Registries (SPRs), we utilized the Ecosyste.ms software package dataset for traditional packages, following the approach outlined in prior work (Schorlemmer et al., 2024). This dataset provides a set of free and open resources for those working to sustain and secure open source software. Ecosyste.ms publishes open data and APIs that map software interdependencies, as well as providing data on the usage, creation, and potential impact of packages. In this work, we used the version of the dataset from October 2023, as this is the closest in time to when the PeaTMOSS dataset was created. This includes the usage of data from two software package registries: PyPI and NPM.
5.3. $C_{1}$: Transformers increases the accessibility of PTM creation and downstream reuse.
@@
5.3.1. Methodology.
We present the metric we developed for claim 1. Metric 1: Preservation rate of libraries to descendents: Some models support multiple libraries. If the claim holds, then descendants of those models should make use of (and thus support) the more reuse-friendly libraries. We examine each library $L$ in turn to assess how frequently its descendants continue to use it. We consider: the library $L$ being assessed, the set of base models $B$ that use library $L$ and at least one other library, the set $D_{b}$ of direct descendants of base model $b\in B$, and a function $S(d,L)$ that returns 1 if descendant $d\in D_{b}$ supports library $L$, otherwise 0. We can then calculate the preservation rate $P_{L}$ of a library $L$ as:
@@
The numerator is the total descendants supporting library $L$, calculated as: $\sum_{b\in B}\sum_{d\in D_{b}}S(d,L)$. The denominator is the total count of direct descendants across all base models in set $B$: $\sum_{b\in B}|D_{b}|$. This equation will give the percentage of direct descendants that support library $L$. The raw count of descendants supporting the library can be found using the numerator.
5.3.2. Results.
Metric 1: Preservation rate of libraries to descendents: As depicted in Figure 5, the Transformers library has the highest survival rate from a parent model to its descendant among all libraries used. Over 80% of descendant models continue to employ the Transformers library, establishing it as the preferred choice for generating descendant models. Contrarily, PyTorch was not favored among PTM descendants, a departure from claims within our systematic literature review that both Transformers and PyTorch are dominant on Hugging Face (Castaño et al., 2024).
Figure 5. The usage proportion of the top-10 libraries that PTMs utilizing at least two different libraries use on Hugging Face. For PTM packages that leverage support at least two libraries, most packages support the transformers library, followed by the Hugging Face promoted SafeTensors library. Most other libraries have little usage in comparison. In contrast to previous work (Castaño et al., 2024), PyTorch is not one of the most popular library to be supported when a PTM package supports more than one library.
Instead, our findings highlight the prevalence of the Transformers and SafeTensors libraries. This suggests a community shift towards prioritizing security, as SafeTensors is designed to replace the commonly used Python pickle library with a more secure container for deploying models. Our results suggest that PTM synthesizers are willing to compromise on functionality and portability to ensure the distribution of more secure PTM packages. Building upon these observations, our findings reveal an opportunity for PTM platforms to further enhance security features. This could involve implementing a visible security rating system for models and developing compatibility layers between secure libraries and those offering greater functionality or portability, encouraging developers to prioritize secure practices without sacrificing functionality or portability.
@@
5.4.1. Methodology.
We present the metrics developed for claim 2.
Metric 2: Turnover of Top PTMs: Drawing on prior work characterizing the stability of top packages over time (Dey and Mockus, 2018), we measured the top-$K$ turnover for each registry. Let $S_{current}$ be the set of $K$ most popular packages in the current snapshot, and $S_{last}$ be the set of top $K$ packages in the last snapshot. We also consider $S_{history}$, the set of packages in any snapshot before $S_{last}$. We then distinguish three categories for packages in $S_{current}$:
@@
For the measurement, we defined popularity by the number of downloads, and examined the top-1000 packages. We obtained snapshots across four dates for both traditional software and PTMs. Data came from several HuggingFace snapshots (datasets: HF Model Metadata–June 2022, PTMTorrent–May 2023, PeaTMOSS–Oct. 2023).
We took a current snapshot of the top 1000 PTMs directly from Hugging Face (April 2024). We used the Ecosyste.ms dataset (NPM, PyPI) for a comparison with traditional software package registries.
@@
5.4.2. Results.
Metric 2: Turnover of Traditional Software and PTM registries: Figure 6 shows the results for Hugging Face and NPM for comparison. The Hugging Face data does not match our interpretation of the claim, with about half of the top-1K Hugging Face PTMs turning over in each snapshot. This high turnover rate suggests that packages on Hugging Face have a shorter lifespan, indicating a dynamic PTM environment where the requirements and preferred models frequently change (Gopalakrishna et al., 2023). In contrast, the traditional registries of PyPI and NPM showed stability with their most popular packages. While 2535 distinct packages rotated through HuggingFace’s top-1K, only 1127 featured for NPM.
@@
The rapid turnover of top models on Hugging Face, in contrast to the stability seen in traditional software packages, reveals that PTM needs are constantly evolving, unlike the more established requirements in conventional software domains. This dynamic environment sees newer models supplant older ones, driven by innovation and rapid adaptation to shifting demands. Further investigation could analyze the lifecycle of these PTMs to determine whether they are newer models briefly appearing or established ones losing prominence, as well as identifying the traits of PTMs or maintainers who consistently stay within the top rankings. These findings have significant implications for improving PTM reuse platforms. Hugging Face and similar platforms should implement robust version control and model lineage tracking systems, enabling users to navigate the rapidly changing landscape of popular models. Users need tools to easily track model evolution, understand version relationships, and identify relevant updates for their specific use cases. Such enhancements would empower users to make informed decisions about model selection and upgrades, minimizing their technical lag in AI advancements.
Metric 3: Number of Descendents of Top Packages: Figure 7 shows the results for this metric. We observe a weak positive correlation between popularity (downloads) and the number of descendent models. While less significant than Metric 2’s findings, our results highlight the need for specialized tools in PTM reuse and guide future research. A more sophisticated recommendation system for PTM platforms is crucial, especially for identifying models suitable for further development. Popularity alone is insufficient for engineers creating new models; more nuanced criteria are necessary. Further analysis should consider factors like model architecture, application diversity, descendant generation patterns, and adaptation potential. This approach could inform the development of informative recommendation systems that help users discover highly adaptable models across various domains and tasks, including those that might otherwise be overlooked due to lower popularity.